The weak acid HIO has a Ka of 2.0×10−11. If a 1.7 M solution of the acid is prepared, what is the pH of the solution? The equilibrium expression is:

HIO(aq)+H2O(l)⇋H3O+(aq)+IO−(aq)

  • Report your answer with two significant figures.
Answer #1

se H10 (ac) + 10 (e) = 10*(as) + Lolac) 0 0 tn th (1.- a a c -a non = 2.0x10 [heo) (1) art 2.0 x10 m - 3:4 X1000 az 5.83x166